For a small manufacturing firm, the unit cost C(x) in dollars of producing x units per day is given by
C(x) = x^2 − 80x + 6000.
How many items should be produced per day to minimize the unit cost?

What is the minimum unit cost?

Answer by lwsshak3(11628)

For a small manufacturing firm, the unit cost C(x) in dollars of producing x units per day is given by
C(x) = x^2 − 80x + 6000.
How many items should be produced per day to minimize the unit cost?
What is the minimum unit cost?
C(x) = x^2 − 80x + 6000.
complete the square:
C(x) = (x^2−80x+1600)-1600 + 6000.
C(x)=(x-40)^2+4400
How many items should be produced per day to minimize the unit cost?40
What is the minimum unit cost?4400 dollars